A total of 640 tickets were sold in play. They were adult or student tickets. There were 60 fewer students tickets sold than adults. How many adult tickets were sold
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52781)

.

a + (a-60) = 640


2a = 640 + 60


2a = 700


a  = 700/2 = 350  adult tickets.     ANSWER


Answer by greenestamps(13200)



Doing the same thing as the formal algebra, but informally....

(1) Add 60 more student tickets. That makes equal numbers of adult and student tickets, and it makes the total number of tickets 700.
(2) So half of the 700 tickets are adult tickets.

ANSWER: 350 adult tickets